# Matching-Based Policy Learning

### v1

@@ 4.2 Compared Methods and Implementation Details
The proposed method is compared with: (a) Q-learning (Qian and Murphy, 2011); (b) Augmented outcome weighted learning (AO-learning) (Liu et al., 2018), which incorporates O-learning with an argumentation term for reducing variability; (c) M-learning (Wu et al., 2020), an alternative policy learning method based on matching; and (d) Policy Tree (Athey and Wager, 2021; Sverdrup et al., 2020), an AIPWE-based policy learning method with fixed-depth decision trees.
@@
For Policy Tree, we also derive the optimal policy within a class of depth-2 decision trees using R package policytree, where the nuisance functions are estimated by R function double_robust_scores in R package grf.
@@
For AO-learning, the parametric and non-parametric variants are also considered, represented by the suffixes -Para and -RF respectively. We estimate the propensity score using logistic regression for $W$ on $(1,X)$ using R function glm in the parametric variant, while regression_forest is used for the non-parametric counterpart. The optimal policy is obtained using R package WeightSVM with Gaussian kernel, with the tuning parameters chosen via 5-fold cross-validation. In addition, for the augmentation term defined in Liu et al. (2018), we fit $Y$ on $X$ employing regression_forest.
For M-learning, we establish the matched sets defined in Wu et al. (2020) using Matching and implement the optimization by WeightSVM, with the kernel and tuning parameters chosen following the same configuration as AO-learning. We employ 1-NN matching, following the argument in the simulation studies of Wu et al. (2020). Although the general $M$-NN matching could be considered, our preliminary simulation studies indicate that the simplest 1-NN matching always yields better performance with less computational burden.
For each scenario, we vary the size of training dataset $n\in\{200,500,1000\}$. The criterion to assess the performance is the empirical value function $\mathbb{P}_{n}[Y(\pi(X))]$, where $\mathbb{P}_{n}$ is the empirical average over an independent testing dataset with size $20000$. Each simulation setup is replicated 200 times.
@@ 4.3 Simulation Results
Figure 1: Boxplot of empirical value functions in different scenarios, where the main effect model is linear and the contrast function is tree. The global optimal value function is 1.25.
@@
We present the simulation results in this subsection. Figure 1 illustrates the performance in scenarios where the main effect model is linear and the contrast function is tree. The global optimal value function $V(\pi^{*}_{opt})$ is 1.25 in this setup. All the approaches exhibit more satisfactory performance with a larger sample size. As expected, the six variants of MB-learning are competitive since the policy $\pi^{*}_{opt}$ is included in the class $\Pi$. Among them, MB-LR-M5 and MB-LASSO-M5 effectively accommodate the main effect model and thus yield the similarly best performance, demonstrating the necessity of bias correction. The results indicate that MB-learning with a larger number of matches $M$ performs slightly better. However, preliminary simulations suggest that MB-learning is insensitive to $M$. Notably, Policy Tree yields the best performance among the remaining methods. Nevertheless, while Policy Tree behaves comparably as MB-learning in Scenarios 1 and 2, it deteriorates in Scenarios 3-5 due to the extreme propensity scores that introduce significant variability. These phenomena align with previous discussions in the literature regarding approaches based on AIPWE (e.g., Kang and Schafer, 2007; Zhao et al., 2019). In contrast, the variants of MB-learning mitigate this issue by circumventing the use of propensity scores and thus exhibit appealing robustness.
@@
Figure 2: Boxplot of empirical value functions in different scenarios, where the main effect model is linear and the contrast function is non-tree. The global optimal value function is 1.36.
@@
Figure 2 presents the results with linear main effect model and non-tree contrast function, where the global optimal value function is 1.36. These scenarios explore the impact of misspecifying the restricted policy class. In general, MB-learning with bias correction and Q-learning-Para outperform the remaining methods. Among them, Q-learning-Para dominates the others in Scenarios 1 and 2, which can be explained by the reasonable accommodation for the outcome mean model $\mu(x,w)$. However, in Scenarios 3 and 5, the variants of MB-learning with bias correction exhibit similarly superior performance, while Q-learning-Para deteriorates in these situations due to the extreme proportion of two treatment groups, particularly with a small sample size ($n=200$). These results demonstrate the competitive performance of MB-learning, which can be attributed to the desirable approximation of fixed-depth decision tree to the global optimal policy $\pi^{*}_{opt}$ and the robustness of matching. In terms of (A)IPWE-based approaches, Policy Tree and AO-learning result in favorable performance in Scenarios 1 and 2, while behaving unsatisfactorily in the remaining settings due to the extreme weights. It is noteworthy that M-learning yields similarly unfavorable performance as MB-learning without bias correction, which can partly be explained by the matching bias.
@@
Figures S1 and S2 in Supplementary Materials display the results with nonlinear main effect model, where the contrast function is tree and non-tree respectively. Generally, MB-LASSO has the best performance. In contrast, MB-LR behaves similarly to MB-LASSO in Scenarios 1, 2, and 4, while resulting in relatively poor performance in Scenarios 3 and 5. This is not surprising. It seems that the conditional bias is relatively minor in the former scenarios, as illustrated by the results of MB-learning without bias correction, which becomes more significant in the latter situations. Hence, bias correction based on linear regression may be insufficient to mitigate the severe conditional bias in Scenarios 3 and 5. In addition, Q-learning-RF is competitive in these settings, outperforming its parametric counterpart as expected. The remaining approaches exhibit performance similar to those in Figures 1 and 2. Notably, a comparison of the results from Figures 2 and S2 indicates that Q-learning-Para outperforms AO-learning-Para when the main effect model is linear, despite the correct specification of logistic regression for propensity score in AO-learning-Para (Figure 2, Scenario 1). However, AO-learning-Para performs slightly better in general when the main effect is nonlinear, including the situations with incorrectly imposed propensity score model (Figure S2, Scenarios 2 and 5, $n=1000$). One possible explanation for these phenomena is that the performance of these two approaches depends on the extent of model misspecification in outcome mean model and propensity score model, which aligns with the findings discussed in Zhao et al. (2019).
@@
To further illustrate the promising performance of MB-learning, we focus on the comparison with Policy Tree with $n$ varies from 200 to 8000. The results are displayed in Figures S3 and S4 in Supplementary Materials. MB-learning with bias correction outperforms Policy Tree by a large margin when the sample size is small, whereas the performance becomes nearly identical with a larger sample size. This phenomenon can be attributed to the cross-fitting technique employed in Policy Tree, where the fit on nuisance functions by random forest may be inadequate with a small sample size. Similar observations have been mentioned in Wu et al. (2022). Furthermore, this finding provides evidence to support the result of convergence rate discussed in Section 3.
5 Application: Treatment Allocation on NSW Program
@@
The NSW Program aimed to provide work experience for job-seekers facing economic and social difficulties before enrollment. The individuals were randomly assigned to either job training program exposed group or unexposed group. The causal effect of the training program on post-treatment earnings has been analyzed in many literature (e.g., LaLonde, 1986; Dehejia and Wahba, 1999, 2002; Abadie and Imbens, 2011; Imai and Ratkovic, 2014). We utilize the NSW datasets, available at https://users.nber.org/ rdehejia/nswdata2.html, to demonstrate the performance of MB-learning in identifying the optimal policy.
@@
We first implement the proposed method on the dataset used in Dehejia and Wahba (1999) (DW), which is a randomized experiment including 185 training program exposed ($W=1$) and 260 unexposed ($W=0$). The 8-dimensional pre-treatment covariates $X$ are considered, including age, education, black (1 if black, 0 otherwise), Hispanic (1 if Hispanic, 0 otherwise), married (1 if married, 0 otherwise), no degree (1 if no degree, 0 otherwise), earnings in 1974 (RE74), and earnings in 1975 (RE75). The outcome of interest $Y$ is earnings in 1978 (RE78). To assess the covariates balance between the groups, we calculate the normalized difference (Abadie and Imbens, 2011), defined as $(\bar{X}_{1}-\bar{X}_{0})/\sqrt{(S^{2}_{1}+S^{2}_{0})/2}$, where $\bar{X}_{w}=\sum_{i:W_{i}=w}X_{i}/n_{w}$, $S^{2}_{w}=\sum_{i:W_{i}=w}(X_{i}-\bar{X}_{w})^{2}/(n_{w}-1)$, and $n_{w}$ represents the number of units in group $W=w$. The normalized difference values are displayed in Table 1, indicating no significant differences in covariates distributions between the groups. We proceed to derive the optimal policy $\hat{\pi}$ that assigns personalized treatments to the job-seekers for maximizing the post-treatment earnings. Due to legal constraints regarding the use of ethnicity for treatment allocation, black and Hispanic are treated as confounders and are thus excluded from identifying the optimal policy.
We apply the six variants of MB-learning discussed in Section 4 to DW dataset and focus on the comparison with Policy Tree, where the optimal policies are estimated within depth-2 decision trees. The estimated policy $\hat{\pi}$ is evaluated by an AIPWE of the value function $V(\hat{\pi})$:
@@
where $\tilde{e}(w,x)$ and $\tilde{\mu}(x,w)$ represent the estimators of $e(w,x)=\Pr(W=w|X=x)$ and $\mu(x,w)$ respectively. Notably, since this dataset is collected from a randomized experiment, $\tilde{e}(w,x)$ is attained by the empirical proportion of the exposed or unexposed group, while $\tilde{\mu}(x,w)$ is derived by random forest. Following the argument in Zhao et al. (2014), we leverage a cross-validated analysis to assess the performance of the estimated policy. Specifically, the dataset is randomly split into five parts with equal sample sizes, with four parts used for estimating the optimal policy and the remaining fifth part used for evaluating the performance. The training and testing sets are permuted five times such that each part is used as the testing dataset in turn. Cross-validated values for each approach are obtained by averaging the results across the five testing sets. This procedure is repeated 100 times. In addition, to further highlight the improvement of personalized treatments, we also explore the one-size-fits-all policy that all units are either exposed or unexposed, where we calculate the AIPWE-adjusted mean.
@@
Table 2 reports the cross-validated results of the value functions, where higher values are desirable. The results indicate that always being exposed is the most beneficial policy for the job-seekers. One possible explanation is that the job training program generally increases earnings for this population, which implies limited heterogeneity. This phenomenon has also been discussed in Zhou et al. (2023). The results can also be attributed to the model misspecification of the restricted policy class, which may lead to the deterioration of MB-learning and Policy Tree, as discussed in Section 4. Among the personalized policies, MB-LASSO-M5 exhibits the best performance, while the other variants of MB-learning are competitive. In contrast, Policy Tree performs the worst in terms of the cross-validated values, which can partly be explained by the unsatisfactory cross-fitting performance in a relatively small sample. Additionally, the optimal policies estimated by MB-LASSO-M5 and Policy Tree are provided in Figure S5 in Supplementary Materials.
@@
Moreover, we investigate the performance of MB-learning on observational data. The DW dataset is composited with additional 429 unexposed units from Current Population Survey (CPS3), resulting in a synthetic observational study (DW-CPS3). The normalized difference values presented in Table 1 suggest that the discrepancy for covariates between the exposed and unexposed group increases, particularly in race, marital status, and earnings. We conduct cross-validated analysis in this observational study, where both $\tilde{e}(w,x)$ and $\tilde{\mu}(x,w)$ are attained by random forest. As shown in Table 2, the variants of MB-learning have superior performance, with MB-M5 yielding the highest cross-validated value and the smallest standard deviation. It is noteworthy that MB-learning slightly deteriorates after bias correction. One possible reason is the unsatisfactory fit for $\mu(x,w)$, which could be highly nonlinear in this population. In addition, the results illustrate that the individualized policies learned by MB-learning outperform the universal policy of always being exposed. These results are likely to be related to the presence of heterogeneity in this composite dataset, given that CPS3 samples are nationally representative, collected across the United States (Smith and Todd, 2005). The job training program may not universally benefit these composite samples, as they were drawn from different local labor markets. On the other hand, Policy Tree is dominated by the variants of MB-learning, exhibiting the highest variability and yielding a cross-validated value worse than always being exposed. We also present the optimal policies learned by MB-M5 and Policy Tree in Figure S6 in Supplementary Materials.
@@
Table 2: Mean cross-validated value functions in different datasets. The standard deviation is in the parenthesis. In DW dataset, the AIPWEs of the value functions relative to “always being exposed" and “always being unexposed" is 6244.9 and 4617.5, respectively. In DW-CPS3 dataset, the AIPWEs of the value functions relative to “always being exposed" and “always being unexposed" is 6614.2 and 5781.6, respectively.

### v2

@@ 4.2 Compared Methods and Implementation Details
The proposed method is compared with: (a) Q-learning (Qian and Murphy, 2011); (b) Augmented outcome weighted learning (AO-learning) (Liu et al., 2018), which improves O-learning with an argumentation term; (c) M-learning (Wu et al., 2020), an alternative policy learning method based on matching; and (d) AIPW-based policy learning with cross-fitting implemented by policy tree (denoted as Policy Tree) (Athey and Wager, 2021; Sverdrup et al., 2020).
@@
For Policy Tree, we derive the optimal policy within a class of depth-2 decision trees using the R package policytree, where the nuisance functions are estimated by the R function double_robust_scores in the R package grf.
@@
For AO-learning, the parametric and non-parametric variants are also considered, represented by the suffixes -Para and -RF respectively. We estimate the propensity score using logistic regression for $W$ on $(1,X)$ using the R function glm in the parametric variant, while regression_forest is used for the non-parametric counterpart. The optimal policy is obtained using the R package WeightSVM with Gaussian kernel where the tuning parameters are chosen via five-fold cross-validation. In addition, for the augmentation term defined in Liu et al. (2018), we fit $Y$ on $X$ employing regression_forest.
For M-learning, we construct the matched sets defined in Wu et al. (2020) using Matching and learn the optimal policy using WeightSVM, where the kernel and tuning parameters are chosen using the same configuration as AO-learning. We only consider 1-NN matching (NN-matching with $M=1$), because our preliminary simulation studies and the numerical results in Wu et al. (2020) indicate that the simplest 1-NN matching yields better performance with less computational burden compared to $M$-NN matching with $M>1$.
For each scenario, we vary the size of training dataset $n\in\{200,500,1000\}$. We use the empirical value function $\mathbb{P}_{n}[Y(\pi(X))]$ as the performance criterion, where $\mathbb{P}_{n}$ is the empirical average over an independent testing dataset with size $20000$. Each simulation setup is replicated 200 times.
@@ 4.3 Simulation Results
Figure 1: Boxplot of empirical value functions where the main effect model is linear and the contrast function is tree. The global optimal value function is 1.25. The three rows represent different sample sizes $n=200,500$, or 1000. The five columns represent different propensity score models presented in Table S1 of Supplementary Material.
@@
We present the simulation results in this subsection. Figure 1 illustrates the performance in five propensity score scenarios with three different sample sizes, where the main effect model is linear and the contrast function is tree. The global optimal value function $V(\pi^{*}_{opt})$ is 1.25 in this setup. All the approaches exhibit more satisfactory performance with a larger sample size. As expected, the six variants of MB-learning are competitive since the policy $\pi^{*}_{opt}$ is included in the class $\Pi$. Among them, MB-LR-M5 and MB-LASSO-M5 correctly fit the main effect model and yield the best performance, demonstrating the necessity of bias correction. The results indicate that MB-learning with a larger number of matches $M$ performs slightly better. However, preliminary simulations suggest that MB-learning is insensitive to $M>5$. Notably, Policy Tree yields the best performance among the remaining methods. Nevertheless, although Policy Tree behaves comparably as MB-learning in Scenarios 1 and 2, its performance deteriorates in Scenarios 3-5 with significant variability due to that the two treatment groups are extremely imbalanced. These phenomena align with discussions in the literature on AIPW-based approaches (e.g., Kang and Schafer, 2007; Zhao et al., 2019). In contrast, the variants of MB-learning mitigate this issue by circumventing the use of propensity score in estimation and learning and exhibit appealing robustness.
@@
Figure 2: Boxplot of empirical value function where the main effect model is linear and the contrast function is non-tree. The global optimal value function is 1.36. The three rows represent different sample sizes $n=200,500$, or 1000. The five columns represent different propensity score models presented in Table S1 of Supplementary Material.
@@
Figure 2 presents the results with linear main effect model and non-tree contrast function, where the global optimal value function is 1.36. These scenarios explore the impact of misspecifying the policy class. In general, MB-learning with bias correction and Q-learning-Para outperform the remaining methods. Among them, Q-learning-Para dominates the others in Scenarios 1 and 2, which can be explained by the reasonable accommodation of the outcome mean model $\mu(x,w)$. However, in Scenarios 3 and 5, the variants of MB-learning with bias correction exhibit superior performance, while Q-learning-Para deteriorates in these situations due to the extreme proportion of the two treatment groups particularly when the sample size is small ($n=200$). These results demonstrate the competitive performance of MB-learning, which can be attributed to the desirable approximation of fixed-depth decision tree to the global optimal policy $\pi^{*}_{opt}$ and the robustness of matching. Among (A)IPW-based approaches, Policy Tree and AO-learning have favorable performance in Scenarios 1 and 2, while they behave unsatisfactorily in the remaining settings. It is noteworthy that M-learning yields similarly unfavorable performance as MB-learning without bias correction, which may be partly explained by the matching bias.
@@
Figures S1 and S2 in Supplementary Material display the results with nonlinear main effect model where the contrast function is tree or non-tree. Generally, MB-LASSO has the best performance. In contrast, MB-LR behaves similarly to MB-LASSO in Scenarios 1, 2, and 4, whereas the performance is relatively poor in Scenarios 3 and 5. This is not surprising because that the conditional bias is relatively minor in the former scenarios, as illustrated by the results of MB-learning without bias correction. The conditional bias becomes more significant in the latter situations. Hence, bias correction based on linear regression may be insufficient to eliminate bias in Scenarios 3 and 5. In addition, Q-learning-RF is competitive in these settings, outperforming its parametric counterpart as is expected. The remaining approaches exhibit similar performance as those in Figures 1 and 2.
@@
We conduct more numerical studies comparing MB-learning with Policy Tree when $n$ varies from 200 to 8000. The results are displayed in Figures S3 and S4 in Supplementary Material. MB-learning with bias correction outperforms Policy Tree by a large margin when the sample size is small, whereas the performance becomes nearly the same when the sample size increases to 8000. This empirical evidence agrees with the theoretical result of convergence rate in Corollary 2. This phenomenon can be attributed to the cross-fitting technique employed in Policy Tree: the nuisance functions are fitted with random forests, which may be inadequate with a small sample size. Similar observations were made in Wu et al. (2022). Finally, Policy Tree encounters the “phase transition” phenomenon in Scenarios 3 and 5, whereas MB-learning does not have this issue in the simulations.
5 Application: Treatment Allocation in the NSW Training Program
@@
The NSW Program aimed to provide work experience for job-seekers encountering economic and social difficulties before enrollment. The individuals were randomly assigned to the training program-exposed group or the unexposed group. The causal effect of the training program on post-treatment earnings has been analyzed in many literature (e.g., LaLonde, 1986; Dehejia and Wahba, 1999, 2002; Abadie and Imbens, 2011; Imai and Ratkovic, 2014). We utilize the NSW datasets, available at https://users.nber.org/ rdehejia/nswdata2.html, to demonstrate the performance of MB-learning in identifying the optimal policy.
@@
We first implement the proposed method using the dataset used in Dehejia and Wahba (1999) (DW), which is a randomized experiment including 185 exposed subjects ($W=1$) and 260 unexposed subjects ($W=0$). The eight-dimensional pre-treatment covariates $X$ are considered as confounders, which consist of age, education, race (1 if black, 0 otherwise), ethnicity (1 if Hispanic, 0 otherwise), marital status (1 if married, 0 otherwise), academic degree (1 if no degree, 0 otherwise), earnings in 1974 (RE74), and earnings in 1975 (RE75). The outcome of interest $Y$ is the earnings in 1978 (RE78). To assess covariate balance between the treatment groups, we calculate the normalized difference (Abadie and Imbens, 2011), defined as $(\bar{X}_{1}-\bar{X}_{0})/\sqrt{(S^{2}_{1}+S^{2}_{0})/2}$, where $\bar{X}_{w}=\sum_{i:W_{i}=w}X_{i}/n_{w}$, $S^{2}_{w}=\sum_{i:W_{i}=w}(X_{i}-\bar{X}_{w})^{2}/(n_{w}-1)$, and $n_{w}$ is the number of units in group $w$, $w=0,1$. The normalized differences are displayed in Table 1, indicating no significant difference in covariate distributions between the groups.
We proceed to derive the optimal policy $\hat{\pi}$ that assigns personalized treatments to the job-seekers with the goal of maximizing the post-treatment earnings. Due to ethical constraints, the black and Hispanic variables are not considered in learning the optimal policy, although they are used in estimating the advantage function or the value function.
We apply the six variants of MB-learning discussed in Section 4 to the DW dataset and focus on the comparison with Policy Tree in Athey and Wager (2021), where the optimal policies by MB-learning and Policy Tree are both estimated within depth-2 decision trees using the R package policytree. The estimated policy $\hat{\pi}$ is evaluated by an AIPW estimator of the value function $V(\hat{\pi})$:
@@
where $\tilde{e}(w,x)$ and $\tilde{\mu}(x,w)$ represent the estimators of $e(w,x)=\Pr(W=w|X=x)$ and $\mu(x,w)$ respectively. Notably, since this dataset is collected from a randomized experiment, $\tilde{e}(w,x)$ is attained by the empirical proportion of the exposed or unexposed group. The $\tilde{\mu}(x,w)$ is derived by random forest. Analogous to the arguments in Zhao et al. (2014), we conduct a cross-validated analysis to assess the performance of the estimated policy. Specifically, the dataset is randomly split into five parts with equal sample sizes, where four parts are used to estimate the optimal policy and the remaining part is used to evaluate the performance. The training and testing sets are permuted five times such that each part is used as the testing dataset in turn. Cross-validated values for each approach are obtained by averaging the results across the five testing sets. This procedure is repeated 100 times. To evaluate the effectiveness of individualized policy assignments, we also explore the one-size-fits-all policy that all units are either exposed or unexposed, where we calculate the AIPW-based adjusted mean value.
@@
Table 2 reports the cross-validated results of the value functions. The results indicate that the universal policy of always being exposed is more beneficial for job-seekers compared to MB-learning and Policy Tree. One possible explanation is that the job training program generally increases earnings for the population, which implies limited heterogeneity. This phenomenon was discussed in Zhou et al. (2023). The results may be partly attributed to misspecification of the restricted policy class, which lead to deteriorated performance of MB-learning and Policy Tree as discussed in Section 4. Among the personalized policies, MB-LASSO-M5 exhibits the best performance, while the other variants of MB-learning are competitive. In contrast, Policy Tree performs the worst in terms of the cross-validated values, which can be explained by the usage of machine learning in estimating nuisance functions with a relatively small sample. The tree structures of the estimated optimal policies by MB-LASSO-M5 and Policy Tree are drawn in Figure S5 in Supplementary Material.
@@
Moreover, we investigate the performance of MB-learning on observational data. The DW dataset is composited with additional 429 unexposed units from Current Population Survey (CPS3), resulting in a synthetic observational study (DW-CPS3). The normalized difference values of the DW-CPS3 dataset presented in Table 1 suggest that the covariate discrepancy between the exposed and unexposed groups is not ignorable particularly for race, marital status, and earnings. We conduct a cross-validated analysis in this observational study, where both $\tilde{e}(w,x)$ and $\tilde{\mu}(x,w)$ are estimated using random forest. As shown in Table 2, the variants of MB-learning have superior performance. Among them, MB-M5 yields the highest cross-validated value and the smallest standard deviation. It is noteworthy that MB-learning slightly deteriorates after bias correction. One possible reason is the unsatisfactory fit of $\mu(x,w)$, which may be highly nonlinear in this population. In addition, the results illustrate that the individualized policies learned by MB-learning outperform the universal policy of always being exposed. These results are likely to be related to the presence of heterogeneity in this composite dataset, given that the CPS3 samples are nationally representative when they were collected across the United States (Smith and Todd, 2005). The job training program may not universally benefit these composite samples, as they were drawn from different local labor markets. On the other hand, Policy Tree is dominated by the variants of MB-learning, exhibiting the highest variability and yielding a cross-validated value worse than the universal policy of always being exposed. We present the tree-structured optimal policies learned by MB-M5 and Policy Tree in Figure S6 in Supplementary Material.
@@
Table 2: Mean cross-validated values in different datasets. The standard deviation is in the parenthesis. In DW dataset, the AIPW estimators of the value function of the “always being exposed" policy and “always being unexposed" policy are 6244.9 and 4617.5, respectively. In DW-CPS3 dataset, the AIPW estimators of the value function of the “always being exposed" policy and “always being unexposed" policy are 6614.2 and 5781.6, respectively.